# Wrinkling and imaging of thin curved sheets

### v1

@@ Abstract
Thin films or sheets subjected to external forces often undergo mechanical instability, leading to regular patterns of wrinkles, folds, and creases. As can be anticipated from the difficulty of flattening a curved globe, any natural curvature of the sheet will have a strong influence on these instabilities. Here, we develop a non-invasive synthetic schlieren imaging technique to image and reconstruct the surface of wrinkling curved sheets, confined to float on water. Our method circumvents the small-amplitude limit of related imaging techniques, and we demonstrate robust means to estimate the reconstruction accuracy. We then evaluate how the sign and magnitude of Gaussian curvature affects the wrinkling of thin curved sheets, and compare observations of the wrinkle wavelength, amplitude and domain structure with recent theoretical predictions. While generally validating model predictions, we find that the assumption of a conserved surface area during wrinkling should be treated with some care. The control of wrinkling behavior demonstrated here can have application in design of liquid lenses, microfluidics, active textured surfaces, and flexible electronic components.
@@ I Introduction
When a thin sheet is confined beyond some critical point, such as by being compressed along its edges, or as the result of differential growth, it responds by forming wrinkles instead of deforming uniformly [1, 2, 3, 4, 5, 6, 7, 8]. This behavior redistributes stress and relaxes excessive strain energy, through the bending of the sheet. The results include a remarkable range of patterns, such as those shown in Fig. 1. As an example, the relatively simple case of the wrinkling of a flat bi-layer, such as a floating membrane [9, 5] or a coating on a thick substrate [1, 10], is governed by two primary factors. The material’s bending resistance, or stiffness, tends to promote wrinkles with larger wavelengths, since a stiffer material better resists bending over short distances. In addition, a restoring force acts to return the sheet to its unwrinkled state by counteracting the energetic cost of deforming the sheet’s substrate. This force discourages large-amplitude deviations of the film’s shape from its reference configuration. The equilibrium between these influences determines the characteristic size and pattern of the wrinkles that emerge on thin sheets in diverse scenarios [11, 12, 13, 14]. An elegant demonstration of this involves a flexible annular sheet, floating on water, as in Fig. 1(a). By adding surfactant outside the annulus, a difference in surface tension can be set up between its outer and inner perimeter. The film is then effectively placed into compression, which results first in the regular radial wrinkling [15] and then the buckling [16] of the film, beyond some critical differences in the surface tension. An indentation force, such as that produced by the droplet of water in Fig. 1(b), can lead to a similar wrinkling response [4].
@@
Figure 1: Patterns due to mechanical instabilities like wrinkling, buckling and folding arise frequently in nature. Examples include: thin sheets floating on water and forced by (a) surface tension [15, 16] or (b) the weight of a droplet [4]; crinkling of the edges of (c) a leaf [17, 18, 3] or (d) sea slug [3]; (e) creasing observed during the swelling of hydrogel beads [19], reminiscent of cortical folding [20, 21]; and (f) wrinkling of a thin spherical shell floating constrained to float on a water bath [22, 23, 24]. Images are reproduced or adapted from (a) [15], (b) [4], (d) [25] and (f) [23], with permission.
@@
More generally, surfaces need not be flat. For example, the differential, volumetric growth of tissues can affect the morphological stability of their variously curved surfaces or interfaces. As in Fig. 1(c–d), this can be observed in the crinkles on the edges of leaves, flowers, sea slugs, or torn plastic bags [17, 18, 3], as well as in the folds of organs like the brain [20, 21] or intestines [26]. Surface curvature is also an important factor for designing wrinkle patterns and transitions into materials and structures, particularly in curved film/substrate and core-shell geometries [27, 19, 28], like the case shown in Fig. 1(e). The curvature anywhere on a surface can be defined by two principal curvatures, $\kappa_{1}$ and $\kappa_{2}$, which give its maximum and minimum curvatures, with respect to a normal or tangent plane. The Gaussian curvature is their product, $\kappa=\kappa_{1}\kappa_{2}$, while the reciprocals of the principal curvatures define the extreme values of the radius of curvature of circles that fit to the surface at that point, $R_{1}$ and $R_{2}$. Since curvature can vary locally, it provides the opportunity for a high level of control over wrinkle patterns. Thin floating shells are beginning to be used to explore this system in more detail [22, 23, 29, 24], as in Fig. 1(f). The mismatch between the inherent curvature of such a film, and the water surface on which it is supported, gives rise to a geometric constraint that is equivalent to lateral confinement [23, 24]. Here, we use the floating shell geometry to experimentally test the predictions of how curvature affects the wavelength, amplitude, and domain structure of wrinkles in thin curved sheets. Films cut from spherical caps, as in Fig. 2(a) are used for the case of positive curvature, $\kappa>0$, while hyperboloids are used as the basis for surfaces with negative Gaussian curvature, $\kappa<0$, as in Fig. 2(b).
@@
Here we introduce a simple ray tracing based, non-invasive imaging method and use it to map out the surfaces of floating wrinkled films with different inherent curvatures, as demonstrated in Fig. 2(c,d). This reformulation of synthetic Schlieren imaging avoids the need for any small-amplitude approximation in the inferred surface height profile.
To make thin shells with positive and negative Gaussian curvatures, we use hemispherical and saddle-shaped substrates, respectively, with a wide range of curvatures. When these curved shells are floated on a water bath, we measure how the wrinkle wavelength, amplitude and domain structure vary with the sign and magnitude of Gaussian curvature. The wrinkle wavelength is independent of curvature, as predicted from models built around the Euler elastica concept [23], while the wrinkle amplitude scales with $\kappa^{1/2}$. This follows from the idea that the wrinkles take up the excess area of the curved film, compared to its two-dimensional projection [22], although stretching of the film leads to considerable differences in how much excess area is available for the cases of positive and negative curvature, respectively. The wrinkles themselves are divided into domains consisting of parallel features, separated by flat domain boundaries, with the domain shapes closely following recently elucidated rules of how reciprocal force diagrams can be used to analyze curved films [29, 24].
@@ II.1 Thin shells
Polydimethylsiloxane (PDMS) was prepared by mixing Sylgard 184 (Dow) base to curing agent in a 10:1 ratio, and degassing under vacuum. Films were fabricated by drop casting PDMS onto the thermo-formed substrates, adapting methods from Ref. [23]. For this, the films were inverted ($\kappa>0$) or steeply inclined ($\kappa<0$), allowing the excess PDMS to drain and to leave an even layer [34]. Films of thickness of $t\simeq 100\,\mu$m were made by repeatedly coating a substrate, and then curing at 80C for 4 hours. To aid in the removal of the polymer films, a layer of PVA glue was deposited and dried on the substrates, prior to the drop casting. This layer was removed by gentle agitation in water, after the cured film had been peeled off its mold. Once removed, films were trimmed to a square cross section with edge lengths of $W=50$ mm.
@@ II.2 Image acquisition and processing
The floating shells were imaged by projecting patterns of dots through the films, from below, as sketched in Fig. 3(a). As projections, random close-packed patterns [40, 41] were generated by placing dots at random onto a two-dimensional space, while enforcing a minimum separation of 0.36 cm between the dots. A series of 30 such patterns, each with dots at an average density of $\rho=$ 5 cm-2, was used for each experiment. A tablet computer was used as a projecting screen, and an overhead digital camera (Nikon D5100) collected images. Once the bath was set up and leveled, the sequence of images was projected and collected automatically, without moving any components. As shown in Fig. 3(b), the undulating surface of the film distorted the pattern of dots, with the convex ridges acting like diverging lenses, and troughs like converging lenses.
@@
Figure 4: Ray tracing diagram showing how light rays are refracted through a water layer with an uneven upper surface. The apparent position of a dot is displaced by a distance $\delta$, as its image passes through that surface.
@@ III Synthetic schlieren imaging
The experimental setup used in this work defines an inverse optics problem: knowing how the wrinkled surface lenses light, we want to work out its shape. A cross-sectional sketch of this problem, in the form of a ray-tracing diagram, is shown in Fig. 4. In this section we describe our approach to solving this inverse problem and reconstructing the shape of the wrinkled film. After validation of the method, we summarize how the films are analyzed to extract the wavelengths and amplitudes of the resulting patterns.
@@ III.1 Surface reconstruction
Consider a flat water bath of initial height $h_{0}$, with a set of point-like dots displayed at its base. The addition of a thin film, of nominal thickness, causes the surface to adopt a deformed shape, with a new height $h$ that depends on position $x$. As viewed from above, a dot initially at position $x_{i}$ will now be seen at an apparent position $x_{f}$, displaced by a distance $\delta=x_{f}-x_{i}$. The ray traced from an overhead camera to this apparent position is refracted by an angle $\theta$ at the air-water interface. In the paraxial approximation (parallel rays, or a distant camera), the deformed height of the water layer at $x_{i}$ is
@@
where $\eta=h^{2}(n-1)/(2n)$. This change of variables leads to a linear system of equations that can be solved without assuming that surface deformations are small in amplitude, i.e. avoiding the constraint that $|h-h_{0}|\ll h_{0}$. Given appropriate boundary conditions, along with the locations and displacements of a large set of discrete points, Eq. 6 can be efficiently and accurately solved for $\eta$, through use of an interpolating function for $\bm{\nabla\cdot\delta}$.
Here, we used the scatteredInterpolant function in Matlab to interpolate smoothly between the displacement data of the imaged dot locations. Taking $n=1.33$ for the water bath, and in the case where the water surface retains its initial height of $h_{0}=5\,$ mm far from the film, we used solvepde to solve the resulting inverse problem for $\eta$ and hence for the height, $h(\mathbf{x})$, of the distorted film surface.
@@ III.2 Validation and feature extraction
To validate and to test the resolving power of our methods, we used artificial input data, corresponding to the exact displacements expected for $N$ points randomly sampling a surface with height $h=h_{0}+A_{\mathrm{p}}\sin{(2\pi x/\lambda)}\sin{(2\pi y/\lambda)}$ ). To ensure the small-angle limit, we used a nearly flat surface with a peak amplitude $A_{\mathrm{p}}=2\,\mu$m. The data for $\bm{\delta}$ were then processed as described in Section III.1, and the height map of the reconstructed surface, $h^{\prime}$, was compared to its expected shape, $h$. As shown in Fig. 5, the relative error of the reconstruction, $\langle|h^{\prime}-h|\rangle/A_{\mathrm{p}}$, was of order one below the Nyquist cutoff, and decreased with a scaling of $1/\rho\lambda^{2}$ as the dot density increased. Our experiments were typically done under conditions where $\rho\lambda^{2}\simeq 200$.
@@
Figure 6: Surface reconstruction and feature detection. (a) The reconstructed surface height map for a typical wrinkled film is shown, with a negative Gaussian curvature defined by radii of curvature $R_{1}=-R_{2}=$ 70 mm. The scale bar is 1 cm. (b) The locations of the maxima (red) and minima (blue) of the height $h$ are shown, superimposed over the film surface. (c) These extrema are determined through cross-sectional profiles, on lines parallel to the film edges, and are used to measure the average wavelength $\lambda$ of the wrinkles.
@@ IV Theory of wrinkling shells
Ratios of the characteristic energies involved in stretching, bending, vertical displacement, and surface tension provide a useful set of dimensionless groups for studying wrinkle patterns. Following the definitions of Tobasco [29], our films have high bendability, $YW^{4}|\kappa|/B\sim 10^{6}$, moderate substrate deformability (or specific thickness [23]), $Y|\kappa|/K_{g}\sim 10$ and weak surface tension (or strong geometric confinement), $YW^{2}|\kappa|/\gamma\sim 10^{3}$. The full range of the parameter space explored in our experiments is given in Table 1, and values for individual films are collated in Table S1 of the online supplemental information [43]. By balancing the bending stiffness $B$ with the substrate stiffness $K_{g}$, the wavelength of any emergent wrinkles is predicted [13, 28, 2] to be

### v2

@@ Abstract
Thin films or sheets subjected to external forces often undergo mechanical instability, leading to regular patterns of wrinkles, folds, and creases. As can be anticipated from the difficulty of flattening a curved globe, any natural curvature of the sheet will have a strong influence on these instabilities. Here, we develop a non-invasive synthetic schlieren imaging technique to image and reconstruct the surface of wrinkling curved sheets, confined to float on water. Our method circumvents the small-amplitude limit of related imaging techniques, and we demonstrate robust means to estimate the reconstruction accuracy. We then evaluate how the sign and magnitude of Gaussian curvature affects the wrinkling of thin curved sheets, and compare observations of the wrinkle wavelength, amplitude and domain structure with recent theoretical predictions. While generally validating model predictions, we find that the assumption of a conserved surface area during wrinkling should be treated with some care. The control of wrinkling behavior demonstrated here can have application in the design of liquid lenses, microfluidics, active textured surfaces, and flexible electronic components.
@@ I Introduction
When a thin sheet is confined beyond some critical point, such as by being compressed along its edges, or as the result of differential growth, it responds by forming wrinkles instead of deforming uniformly [1, 2, 3, 4, 5, 6, 7, 8]. This behavior redistributes stress and relaxes excessive strain energy, through the bending of the sheet. The results include a remarkable range of patterns, such as those shown in Fig. 1. As an example, the relatively simple case of the wrinkling of a flat bi-layer, such as a floating membrane [9, 5] or a coating on a thick substrate [1, 10], is governed by two primary factors. The material’s bending resistance, or stiffness, tends to promote wrinkles with larger wavelengths, since a stiffer material better resists bending over short distances. In addition, a restoring force acts to return the sheet to its unwrinkled state by counteracting the energetic cost of deforming the sheet’s substrate. This force discourages large-amplitude deviations of the film’s shape from its reference configuration. The equilibrium between these influences determines the characteristic size and pattern of the wrinkles that emerge on thin sheets in diverse scenarios [11, 12, 13, 14]. An elegant demonstration of this involves a flexible annular sheet, floating on water, as in Fig. 1(a). By adding surfactant outside the annulus, a difference in surface tension can be set up between its outer and inner perimeters. The film is then effectively placed into compression, which results first in the regular radial wrinkling [15] and then the buckling [16] of the film, beyond some critical differences in the surface tension. By placing a droplet of water on a film, as in Fig. 1(b), the pull provided by the droplet’s surface tension acting on the film, along with the indentation force of its own weight, can lead to a similar wrinkling response [4].
@@
Figure 1: Patterns due to mechanical instabilities like wrinkling, buckling and folding arise frequently in nature. Examples include: thin sheets floating on water and forced by (a) surface tension differences across an annulus [15, 16] or (b) the weight and surface tension of a droplet [4]; crinkling of the edges of (c) a leaf [17, 18, 3] or (d) sea slug [3]; (e) creasing observed during the swelling of hydrogel beads [19], reminiscent of cortical folding [20, 21]; and (f) wrinkling of a thin spherical shell constrained to float on a water bath [22, 23, 24]. Images are reproduced or adapted from (a) [15], (b) [4], (d) [25] and (f) [23], with permission.
@@
More generally, surfaces need not be flat. For example, the differential, volumetric growth of tissues can affect the morphological stability of their variously curved surfaces or interfaces. As in Fig. 1(c–d), this can be observed in the crinkles on the edges of leaves, flowers, sea slugs, or torn plastic bags [17, 18, 3], as well as in the folds of organs like the brain [20, 21] or intestines [26]. Surface curvature is also an important factor for designing wrinkle patterns and transitions into materials and structures, particularly in curved film/substrate and core-shell geometries [27, 19, 28], like the case shown in Fig. 1(e). The curvature anywhere on a surface can be defined by two principal curvatures, $\kappa_{1}$ and $\kappa_{2}$, which give its maximum and minimum curvatures, with respect to a normal or tangent plane. The Gaussian curvature is their product, $\kappa=\kappa_{1}\kappa_{2}$, while the reciprocals of the principal curvatures define the extreme values of the radius of curvature of circles that fit to the surface at that point, $R_{1}$ and $R_{2}$. Since curvature can vary locally, it provides the opportunity for a high level of control over wrinkle patterns.
Thin floating shells are beginning to be used to explore the wrinkling of curved sheets in more detail [22, 23, 29, 24], as in Fig. 1(f). The mismatch between the inherent curvature of such a film, and the fluid surface on which it is supported, gives rise to a geometric constraint that is equivalent to lateral confinement [23, 24]. Surprisingly, even a small amount of curvature can easily dominate over surface tension effects in the level of confinement that it represents (e.g. see Fig. S2 of Ref. [22]). This system has also been shown to be particularly rich in patterns, with studies exploring the effects of the size and shape of a piece cut out of a curved surface on the wrinkle patterns [22, 23, 24], for example. Here, we use the floating shell geometry to experimentally test the predictions of how curvature affects the wavelength, amplitude, and domain structure of wrinkles in thin curved sheets. Films cut from spherical caps, as in Fig. 2(a) are used for the case of positive curvature, $\kappa>0$, while hyperboloids are used as the basis for surfaces with negative Gaussian curvature, $\kappa<0$, as in Fig. 2(b).
@@
A recent review of free-surface imaging [32] gives a detailed comparison of surface reconstruction methods, including synthetic schlieren techniques and other non-intrusive methods like fringe projection profilometry and 3D scanning systems, which can capture surface topographies but which face limitations due to complexity, cost, and adaptability to transparent surfaces or surfaces of low reflectivity. Refraction-based methods, like synthetic schlieren, are noted for their high accuracy in slope measurement, and for simplicity in implementation [32].
Here we introduce a simple ray-tracing based, non-invasive imaging method and use it to map out the surfaces of floating wrinkled films with different inherent curvatures, as demonstrated in Fig. 2(c,d). This reformulation of synthetic schlieren imaging avoids the need for any small-amplitude approximation in the inferred surface height profile.
To make thin shells with positive and negative Gaussian curvatures, we use hemispherical and saddle-shaped substrates, respectively, with a wide range of curvatures. When these curved shells are floated on a water bath, we measure how the wrinkle wavelength, amplitude and domain structure vary with the sign and magnitude of the Gaussian curvature, $\kappa$. The wrinkle wavelength is independent of curvature, as predicted from models built around the Euler elastica concept [23], while the wrinkle amplitude scales with $\kappa^{1/2}$. This follows from the idea that the wrinkles take up the excess area of the curved film, compared to its two-dimensional projection [22], although stretching of the film leads to differences in how much excess area is available for the cases of positive and negative curvature, respectively. The wrinkles themselves are divided into domains consisting of parallel features, separated by domain boundaries, with the domain shapes closely following recently elucidated rules of how reciprocal force diagrams can be used to analyze curved films [29, 24].
@@ II.1 Thin shells
Polydimethylsiloxane (PDMS) was prepared by mixing Sylgard 184 (Dow) base to curing agent in a 10:1 ratio, and degassing under vacuum. Films were made by drop casting PDMS onto the thermo-formed substrates, adapting methods from Ref. [23]. For this, the films were inverted ($\kappa>0$, see methods in Ref. [34]) or steeply inclined ($\kappa<0$), allowing the excess PDMS to drain and leave an even layer. Films of thickness of $t\simeq 100\,\mu$m were made by repeatedly coating a substrate, and then curing at 80C for 4 hours. To aid in the removal of the polymer films, a layer of PVA glue was deposited and dried on the substrates, prior to the drop casting. This layer was removed by gentle agitation in water, after the cured film had been peeled off its mold. Once removed, films were trimmed to a square cross section with edge lengths of $W=50$ mm.
@@ II.2 Image acquisition and processing
The floating shells were imaged by projecting patterns of dots through the films, from below, as sketched in Fig. 3(a). As projections, random close-packed patterns [40, 41] were generated by placing dots at random onto a two-dimensional space, while enforcing a minimum separation of 0.36 cm between the dots. A series of 30 such patterns, each with dots at an average density of $\rho=$ 5 cm-2, was used for each experiment. A tablet computer was used as a projecting screen, and an overhead digital camera (Nikon D5100) collected images. All experiments were done under calm conditions, with no other users in the lab, and using computer controls to trigger the imaging sequence. Otherwise, a ground-floor lab built directly on solid building foundations, with the experimental setup positioned on top of a heavy ceramic bench-top, was found to provide sufficient passive control to minimize surface vibrations of the water bath. Once the bath was set up and leveled, the sequence of images was projected and collected automatically, without moving any components. As shown in Fig. 3(b), the undulating surface of the film distorted the pattern of dots, with the convex ridges acting like diverging lenses, and troughs like converging lenses.
@@
Figure 4: Ray tracing diagrams showing how light rays are refracted through a water layer with an uneven upper surface. (a) The apparent position of a dot is displaced by a distance $\delta$, as its image passes through that surface. (b) Close-up, the film can be considered as a thin surface layer of thickness $t$, with negligible effect on the displacement pattern if thmuch-less-$t\ll h$.
@@ III Synthetic schlieren imaging
The experimental setup used in this work defines an inverse optics problem: knowing how the wrinkled surface lenses light, we want to work out its shape. A cross-sectional sketch of this problem, in the form of a ray-tracing diagram, is shown in Fig. 4(a). In this section we describe our approach to solving this inverse problem and reconstructing the shape of the wrinkled film. After validation of the method, we summarize how the films are analyzed to extract the wavelengths and amplitudes of the resulting patterns.
@@ III.1 Surface reconstruction
Consider a flat water bath of initial height $h_{0}$, with a set of point-like dots displayed at its base. The addition of a thin film, of nominal thickness, causes the surface to adopt a deformed shape, with a new height $h$ that varies with position. As viewed from above, a dot initially at position $x_{i}$ will now be seen at an apparent position $x_{f}$, displaced by a distance $\delta=x_{f}-x_{i}$. The ray traced from an overhead camera to this apparent position is refracted by an angle $\theta$ at the air-water interface. For this, we assume that the surface height is single-valued (i.e. that it doesn’t fold over itself) and that any ray intersects the surface only at a single point. In the paraxial approximation (parallel rays, or a distant camera), the deformed height of the water layer at $x_{i}$ is
@@
where $\eta=h^{2}(n-1)/(2n)$. Given appropriate boundary conditions, along with the locations and displacements of a large set of discrete points, Eq. 6 can be efficiently and accurately solved for $\eta$, through use of an interpolating function for $\bm{\nabla\cdot\delta}$.
Previous approaches to synthetic schlieren imaging have simplified and solved Eq. 4 in the small-amplitude limit [30, 32], by making the approximation that $\bm{\delta}\approx((n-1)/n)h_{0}\bm{\nabla}h$. The change of variables we use to develop Eq. 6 instead leads to a linear system of equations that can be solved without assuming that surface deformations are small in amplitude, i.e. avoiding the constraint that $|h-h_{0}|\ll h_{0}$. It also avoids the over-constrained nature of inverting Eq. 5, and allows access to standard solvers of the Poisson equation.
Here, we used the scatteredInterpolant function in Matlab to interpolate smoothly between the displacement data of the imaged dot locations. Taking $n=1.33$ for the water bath, and in the case where the water surface retains its initial height of $h_{0}=5\,$ mm far from the film, we used solvepde to solve the resulting inverse problem for $\eta$ and hence for the height, $h(\mathbf{x})$, of the distorted film surface. Finally, we note that, as long as the film is thin compared to the water bath, thmuch-less-$t\ll h$, the different refractive index of the film has only a minimal effect on these calculations, and so is ignored here for simplicity. For example, calculating ray deflections using the three-layer model sketched in Fig. 4(b) would lead to corrections of less than $10\,\mu$m in $\delta$, for angles up to $\varphi_{i}=1.0$ and assuming a thickness of $t=100\,\mu$m and refractive index of $n_{f}=1.40$ for the PDMS film.
@@ III.2 Validation and feature extraction
To validate and to test the resolving power of our methods, we used artificial input data, corresponding to the exact displacements expected for $N$ points randomly sampling a surface with height $h=h_{0}+A_{\mathrm{p}}\sin{(2\pi x/\lambda)}\sin{(2\pi y/\lambda)}$ ). To ensure the small-angle limit, we used a nearly flat surface with a peak amplitude of $A_{\mathrm{p}}=2\,\mu$m. The data for $\bm{\delta}$ were then processed as described in Section III.1, and the height map of the reconstructed surface, $h^{\prime}$, was compared to its expected shape, $h$. As shown in Fig. 5, the relative error of the reconstruction, $\langle|h^{\prime}-h|\rangle/A_{\mathrm{p}}$, was of order one below the Nyquist cutoff, and decreased with a scaling of $1/\rho\lambda^{2}$ as the dot density increased. Our experiments were typically done under conditions where $\rho\lambda^{2}\simeq 200$.
@@
Finally, to evaluate the noise levels introduced by any vibrations, we compared two sets of dot patterns taken through a water bath at different times, in the absence of any floating film. The average displacement of these dots showed a random distribution, with an average value of $\langle|\bm{\delta}|\rangle=23\,\mu$m, close to the 19 $\mu$m pixel resolution of the camera used. The surface reconstructed from these displacements was flat, with a root-mean-squared roughness of 29 $\mu$m.
Figure 6: Surface reconstruction and feature detection. (a) The reconstructed surface height map for a typical wrinkled film is shown, with a negative Gaussian curvature defined by the two radii of curvature $R_{1}=-R_{2}=$ 70 mm. The scale bar is 1 cm. (b) The locations of the maxima (red) and minima (blue) of the height $h$ are shown, superimposed over the film surface. (c) These extrema are determined through cross-sectional profiles, on lines parallel to the film edges, and are used to measure the average wavelength $\lambda$ of the wrinkles.
@@ IV Theory of wrinkling shells
Ratios of the characteristic energies involved in stretching, bending, vertical displacement, and surface tension provide a useful set of dimensionless groups for studying wrinkle patterns. Following the definitions of Tobasco [29], our films have high bendability, $YW^{4}|\kappa|/B\sim 10^{6}$, moderate substrate deformability (or specific thickness [23]), $Y|\kappa|/K_{g}\sim 10$ and weak surface tension (or strong geometric confinement), $YW^{2}|\kappa|/\gamma\sim 10^{3}$. The full range of the parameter space explored in our experiments is given in Table 1, and values for individual films are collated in Table S1 of the online supplemental information [43]. In practical terms, this parameter range means that surface tension will be negligible, as has been demonstrated with experiments involving even thinner PDMS shells [22]. By balancing the bending stiffness $B$ with the substrate stiffness $K_{g}$, the wavelength of any emergent wrinkles is predicted [13, 28, 2] to be